OpenStudy (anonymous):

subtract: y/y-2 - 8+y/2-y

OpenStudy (anonymous):

This is tricky... you need to know this fact. 2 - y = -(y - 2) you can always flip numbers around a - sign and multiply by a - so 5 - x = -(x - 5)

OpenStudy (anonymous):

In this case you want the 2 - y to be y - 2 because you need a common denominator. so you turn the bottom in to y - 2 and then you put a - sign out in front... In this case that negative sign out front would turn your problem into a + sign y/y-2 + 8 + y/y - 2 Now you have common denominators y + 8 + y -------- y - 2 2y + 8 ------ y - 2

OpenStudy (anonymous):

omg hi again!!!

OpenStudy (anonymous):

this was our next question

OpenStudy (anonymous):

OK... I just found you... do you get this?

OpenStudy (anonymous):

i dont understand how you get the common denom....

OpenStudy (anonymous):

Well... y - 2 and 2 - y are opposites of each other.... you should be able to see that if you flip the one on the right around the minus sign it becomes the y - 2 ex. 2 - y flips to y - 2 OK?

OpenStudy (anonymous):

then you have to change the subtraction to addition??

OpenStudy (anonymous):

Yes... because it is the opposite you have to put a - sign out in front, and actually you change the sign of the problem... so in this case the subtraction becomes addition. This is a hard concept.

OpenStudy (anonymous):

hmmm then what

OpenStudy (anonymous):

Then you have a common denominator... so if you look above to my steps you will see that you put everything on the top over y - 2 because you now have a common denominator

OpenStudy (anonymous):

7/T-2 - 6/T^2-2T - 3/T is the LCD T^2-2T?

OpenStudy (anonymous):

7 6 3 ---- - ----------- - ------ T - 2 T^2 - 2T T T - 2 factors into (T - 2) T^2 - 2T factors into T(T - 2) T factors into T Common denominator is T(T - 2) or T^2 - 2T easier if you write it T(T - 2)

OpenStudy (anonymous):

T(T - 2) * 7 T(T - 2) * 6 T(T - 2) * 3 ------------ - ---------- - ---------- (T - 2) T(T - 2) T

OpenStudy (anonymous):

The (T - 2)'s cancel in the first fraction The T and (T - 2) cancel in the second fraction The T cancels in the third fraction

OpenStudy (anonymous):

7t-6-3t-2 ??

OpenStudy (anonymous):

7T - 6 - 3(t - 2) 7T - 6 - 3T + 6 4T

OpenStudy (anonymous):

isnt it -6?

OpenStudy (anonymous):

left with 4t-12

OpenStudy (anonymous):

The - sign goes with the 3 so it is actually -3(T - 2) and -3 * -2 = +6

OpenStudy (anonymous):

ohhh got it

OpenStudy (anonymous):

meeee tooooooo!

OpenStudy (anonymous):

YEAH!!!!!

OpenStudy (anonymous):

5-x/x-1 / 2-x/x+1

OpenStudy (anonymous):

5 - x 2 - x ---- divided by ----- x - 1 x + 1 write the divide sign that has a --- with a dot above and below. Are you sure this is what you want.

OpenStudy (anonymous):

noooo wait hold on

OpenStudy (anonymous):

X 5- ---- X-1 ___________ X 2- ---- X+1

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

OK... First we will simplify the top. 5 x - - ---- 1 x- 1 what is the common denominator?

OpenStudy (anonymous):

x-1?

OpenStudy (anonymous):

(x - 1)*5 x ------- - ----- x - 1 x - 1

OpenStudy (anonymous):

5x - 5 x ----- - ----- x - 1 x - 1 4x - 5 ------ x - 1 ok?

OpenStudy (anonymous):

5X+1-X --------- X-1

OpenStudy (anonymous):

GOT IT

OpenStudy (anonymous):

Yes... now add the x's

OpenStudy (anonymous):

Now let's do the same to the bottom.

OpenStudy (anonymous):

X+2 ----- X+1

OpenStudy (anonymous):

2 x - - ---- 1 x + 1 (x + 1)*2 x -------- - ---- (x + 1)*1 x + 1 (x + 1) *2 x -------- - ----- x + 1 x + 1 2x + 2 - x --------- x + 1 x + 2 ---- x + 1

OpenStudy (anonymous):

OOPS... go back to the first one... you have 4x - 1 look at mine... it is 4x - 5.. you have to distribute.

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

4x - 5 x + 2 ------ divided by ------ x - 1 x + 1 Now.... when ever you have a fraction divided by a fraction. you ALWAYS multiply by the reciprocal

OpenStudy (anonymous):

4X+5 ------ X+1 -------- X+2 --- X+1

OpenStudy (anonymous):

4x - 5 x + 1 ------ x ----- x - 1 x + 2

OpenStudy (anonymous):

top right and bottom left cancel out?

OpenStudy (anonymous):

4X+5 ------ X+2

OpenStudy (anonymous):

amazing... go back and look at the numerator.. it is 4x - 5

OpenStudy (anonymous):

4X-5 ------ X+2 ????? I HOPE

OpenStudy (anonymous):

I wish they did cancel... they usually do.. but not here a (x - 1) can ONLY cancel with an (x - 1) and a (x + 1) can ONLY cancel with (x + 1) Usually you CAN cancel, but not in this problem. You would multiply the numerators (FOIL METHOD) then multiply the denominators (FOIL METHOD) and you would have your answer

OpenStudy (anonymous):

so now what!?!?!!? omg im lost now

OpenStudy (anonymous):

(4x - 5)(x + 1) ------------- (x - 1)(x + 2) 4x^2 + 4x - 5x - 5 ---------------- x^2 + 2x - x - 2 4x^2 - x - 5 ----------- x^2 + x - 2

OpenStudy (anonymous):

Unless you did not type in the original problem correctly? otherwise this is the answer.

OpenStudy (anonymous):

i got that

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

i wanna cry this is hard

OpenStudy (anonymous):

solve for X: (4x-7)^2 = -27 :)

OpenStudy (anonymous):

They are just a lot of steps.

OpenStudy (anonymous):

(4x - 7)^2 = -27 Are you sure this is correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yessssssssss

OpenStudy (anonymous):

OK.. so you know what imaginary numbers are i's

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

OK.. what if I gave you this problem... do you know what to do. x^2 = 16

OpenStudy (anonymous):

yes square root both sides

OpenStudy (anonymous):

Good... so in this problem you have (something)^2 = -27 so you square root both sides

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

(4x - 7)^2 = -27 4x - 7 = sqrt(-27)

OpenStudy (anonymous):

sqrt(-27) is -3i sqrt(3) do you understand this?

OpenStudy (anonymous):

\[3i \sqrt{3?}\]

OpenStudy (anonymous):

yess!

OpenStudy (anonymous):

OOPS you are correct is is positive

OpenStudy (anonymous):

4x - 7 = 3i sqrt(3) 4x = 3i sqrt(3) + 7 x = 3i sqrt(3) + 7 ------------ 4

OpenStudy (anonymous):

\[9+3x-2x ^{2} \over x ^{2}-16 \]

OpenStudy (anonymous):

wait im not done hold on!

OpenStudy (anonymous):

all that divided by \[4x ^{3}-9x \over 2x ^{2}+5x-12\]

OpenStudy (anonymous):

simplify and state restricitons

OpenStudy (anonymous):

-2x^2 + 3x + 9 2x^2 + 5x - 12 --------------- x ------------- x^2 - 16 4x^3 - 9x since it is a fraction divide by a fraction.. we are going to turn it into a multiply problem right away by using the reciprocal. OK?

OpenStudy (anonymous):

ok hold on lemme take all that in

OpenStudy (anonymous):

then factor?

OpenStudy (anonymous):

yes we have to factor all four... but it is difficult to factor when there is a negative in front of the x^2 so for the first fraction in the numerator we are going to factor out a - sign first. - (2x^2 - 3x - 9)

OpenStudy (anonymous):

-(2x + 3)(x - 3) (2x - 3)(x + 4) -------------- x ------------- (x + 4)(x - 4) x(4x^2 - 9) x(2x + 3)(2x - 3)

OpenStudy (anonymous):

-(2x+3)(x-3) (x+4) (2x-3) ---------- ------------ (x+4) (x-4) x(2x+3) (2x-3)

OpenStudy (anonymous):

now we can cancel out right?

OpenStudy (anonymous):

Yes... now cancel anything in ( )'s with like ( )'s Anything on the top with anything on the bottom that matches.

OpenStudy (anonymous):

x-3 ----- x(x-4)

OpenStudy (anonymous):

close you still have that - sign on top -(x - 3) ------- x(x - 4) they may want it like -x + 3 ------ x^2 - 4x it is all the same...

OpenStudy (anonymous):

Now you also said you had to tell restrictions.

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

OK... go back to the factored fractions. The rule is you can NEVER divide by 0 so whatever makes those ( )'s 0 would be a restriction.

OpenStudy (anonymous):

We are only concerned with the denominators.

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

(x + 4)(x - 4) x(2x + 3)(2x - 3)

OpenStudy (anonymous):

-4,4 -3/2, 3/2

OpenStudy (anonymous):

And one more you need 0 because of the single x Every where there is an x there is a restriction.

OpenStudy (anonymous):

blexting you are great!!! we have 3 more can we keep going?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

way better then our professor!

OpenStudy (anonymous):

X^2-9 ---------------- X^3+X^2-9X-9 SIMPLIFY

OpenStudy (anonymous):

Factor